OpenStudy (anonymous):

Find an equation of a tangent line to the parabola y=X^2 at the point (1,-1)

OpenStudy (anonymous):

Forgot to add but this is a precalculus problem and using derivatives is not allowed

OpenStudy (callisto):

Can you use limit?

OpenStudy (anonymous):

I'm not really sure my teacher never said anything about limits. I was out of school that day and she never really caught me up

OpenStudy (callisto):

I guess limit is a pre-calculus topic.. Do you know how to find limit?

OpenStudy (anonymous):

Unfortunately no

OpenStudy (callisto):

slope = \(\lim_{h\rightarrow 0}\frac{f(x+h)-f(x)}{h} \) f(x) = y = x^2 f(x+h) = (x+h)^2 Can you expand f(x+h)?

OpenStudy (callisto):

expand f(x+h) = expand (x+h)^2

OpenStudy (loser66):

one question, (1,-1) is not in parabola, how to find the tangent line at this point? or the tangent line of parabola pass through that point?